What is the maximum number of electrons that can be accommodated in each of the following?
i) n = 3
iii) a 4d subshell
iv) a 2p orbital
iv) n = 3, l = 0, ml = 0, and ms = - ½
1
Expert's answer
2020-06-22T08:52:23-0400

i) maximum electrons is given by 2n2

so maximum electrons in n=3 is 2.3.3=18


ii) for a d subshell l=2,

no. of orbitals=2l+1=2*2+1=5

no. of electrons that can be accommodated="2\times5=10"


iii) for a p orbital, l=1

no. of orbitals =2l+1 = 2*1+1=3

no. of electrons that can be accommodated="2\times3=6"


iv) answer = 1